OpenStudy (anonymous):

An urn contains six red balls, five white balls, and four black balls. Four balls are drawn from the urn at random without replacement. For each red ball drawn, you win $10, and for each black ball drawn, you lose $15. Let X represent your net winnings. Compute E(X), your expected net winnings.

OpenStudy (kropot72):

First you need to find the probabilities for 1, 2, 3 and 4 red balls in a draw of four balls, and the probabilities for 1, 2, 3 and 4 black balls in a draw of four balls. \[\large P(1\ red)=\frac{6C1\times 9C3}{15C4}\] \[\large P(2\ red)=\frac{6C2\times 9C2}{15C4}\] \[\large P(3\ red)=\frac{6C3\times 9C1}{15C4}\] \[P(4\ red)=\frac{6C4}{15C4}\]

OpenStudy (anonymous):

what does C represents

OpenStudy (kropot72):

For example, 9C3 means the number of combinations of 9 different things taken 3 at a time. It can also be read as '9 choose 3'.

OpenStudy (kropot72):

Have you studied combinations?

OpenStudy (anonymous):

i haven't studied it yet just studying ahead of class. You getting ready for it. If you could shed some light on it. I would really appreciate it.

OpenStudy (anonymous):

in combination so far i know is that order doesn't matter. n!/r! (n-r)!

OpenStudy (kropot72):

Many calculators have the combinations function. Using a suitable calculator the results for the equations that I posted are: P(1 red) = 0.369; P(2 red) = 0.396; P(3 red) = 0.132; P(4 red) = 0.011.

OpenStudy (kropot72):

Now you need to calculate the probabilities of 1, 2, 3 and 4 black balls in a draw of four balls.

OpenStudy (kropot72):

\[\large P(1\ black)=\frac{4C1\times 11C3}{15C4}\]

OpenStudy (kropot72):

Having found all the required probabilities, each value of red probability is multiplied by $10 and each value of black probability is multiplied by $15 and assigned a negative value. These products are then summed to find the expected value.

OpenStudy (kropot72):

\[\large E(X)=\sum_{}^{}xp(x)\]